A banner in the shape of an isosceles triangle has a base that is 5 inches shorter than either of the equal sides. If the perimeter of the banner is 37 inches, then what is the length of the equal sides?
Answer by checkley77(12844)

B=S-5
2S+B=37
2S+(S-5)=37
3S=37+5
3S=42
S=42/3
S=14 INCHES FOR THE EQUAL SIDES.
B=14-5=9 INCHES FOR THE BASE.
PROOF:
2*14+9=34
28+9=37
37=37